Can a entire population be measured not just a sample?
Yes it can; given that the population number is reasonable and it can be done in a reasonable amount of time.
Why is the jumping off a building is death?
cause you break every bone of your body plus your head
Answer:
Like so many events there is not just one outcome of jumping off a building. Many people die from jumping from tall building but survive jumps from smaller buildings. Human survival of falls from the 17th floor of buildings has been recorded and several parachutists have survived falls from altitudes far taller than any building.
Death in falls usually occurs when the falling person hitting head on when landing or if the body lands on a hard (unyielding) surface. A unyielding surface causes a sudden deceleration. The body acts as fluid and spreads sidewards rupturing many internal organs and causing the brain to splash against the downward side of the skull.
Interestingly a significant amount of work has been done on the defenestration of cats in New York city (Defenestration is the technical term for jumping ot of windows). Cats dropping from 6 floors up or so die. Above the 9th floor they live. This is thought to be due to cats reaching a lower terminal velocity if they have time to relax on the way down.
Probability of event occurring to it not occurring?
p/(1-p)
The sum of the probability of an event occurring plus it not occurring equals 1.
Where did scalping of heads originate?
The Dutch brought scalping practices with them when they colonized. It is a Dutch Tradition to scalp human beings.
How many ways can you select 3 cards from a regular deck of 52 cards?
If it matters in what order you draw them, then the number of permutations of 52 things taken three at a time is 52! - (52-3)! = 52 * 51 * 50 = 132,600.
If it does not matter in what order you draw them, the usual case, then the number of combinations of 52 things taken 3 at a time is (52! - (52-3)!) / 3! = 52 * 51 * 50 / 3 / 2 / 1 = 22,100.
What is the probability of a quarter?
I'm assuming you are asking what is the probability (P) of flipping a quarter.
This answer really depends upon how many times up are going to flip it.
If you are flipping it once, you have a 50% chance that it will land on heads and a 50% chance that it will land on tails. Either way the sum of your probabilities will add up to 1, meaning that there is a 100% chance that something will occur (see probability rules).
EX: Let H= heads and let T=tails
∑P= P(H)+P(T)=0.5+0.5=1
However, let's say you were going to flip a coin 3 times and were wanting to know what the probability of getting at least 1 tail was. You would approach the problem this way:
P( at least 1 tail)=?
Next, you want to find the compliment (the opposite of what you are starting with). So the opposite of getting one tail is getting no tails. This is the same as getting all heads.
P(no tails)=P(all heads)
P( all heads)= P(H)3 Heads is cubed because you are flipping the coin 3
= P(0.5)3 times and want all the outcomes to be heads.
= 1/8
By knowing that the outcome plus its compliment add up to equal 1 you get:
P( all heads) + P( at least 1 tail)=1
P( at least 1 tail) = 1- P( all heads)
P( at least 1 tail) = 1- 1/8
P( at least 1 tail) = 7/8
So the probability of flipping a coin 3 times and getting a least 1 tail is 7/8. In other words, it's very likely that it will land on tails one of those three times.
How many integers between 2000 and 3999 have a ones digit that is a prime number?
There are 800 such numbers.
It is 1/2.
How does probability relate to the punnet square?
A Punnet square is a way of representing the possible combinations of genotypes in an offspring of two parents with differing genes. It is used to determine the probability that the offspring has a particular genotype.
What is the probability of 0.5?
There is insufficient information for us to even begin to understand this question. Please edit the question to include more context or relevant information.
ask from anyone else
What is the probability that the card chosen is a queen?
For a single random choice from a standard deck, the probability is 1/13.
For a single random choice from a standard deck, the probability is 1/13.
For a single random choice from a standard deck, the probability is 1/13.
For a single random choice from a standard deck, the probability is 1/13.
Why you pick 24 as lucky number?
Only because you are superstitious enough to believe that there is anything like a lucky number.
75%
What is the probability of producing dwarf plants with white flowers?
It depends upon what the parents were.
If both parents were dwarf and white, then it is 100%
If both parents were heterozygous for both traits, then it is 1/16.
If either parent were homozygous tall and colored, then it is 0.
Make a Punnett Square to get all of the probabilities if you know the genotypes of the parents.
How many poker hands with 5 cards contain 3 clubs and two hearts?
13 13
# of poker hands with 3 clubs & 2 heards = ( ) ( )
3 2
= (13!)( 13!) / (10!*3!*11!*2)
= (13!)( 13!) / (10!*11!*12)
= (13!)( 13!) / (10!*12!)
= 13 * 13* 12 * 11
= 22308 #